OpenStudy (anonymous):

a metal rod 1.o m long and 10milimmeeter in diameter is drawn out into wire 1.0mm indiameter. a. what is the length of the wire? b. compare the resistance of the rod with that of the wire

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here the volume of the wire has to to be equal to the volume of the metal rod

OpenStudy (anonymous):

yess

OpenStudy (michele_laino):

so, being the volume V of a cylinder given by the subsequent formula: \[V = \frac{{\pi {D^2}}}{4} \times L\] where L is the length and D is the diameter of the cylinder, we can write: \[\Large \frac{{\pi \times {{10}^2}}}{4} \times {10^3} = \frac{{\pi \times {1^2}}}{4} \times L\] where L is the requested length. Please note that I have measured the diameters and the known length of the cylinder in millimeters

OpenStudy (anonymous):

yes then whats nest

OpenStudy (anonymous):

V is not given too

OpenStudy (michele_laino):

we have to compute the length L from that formula above